• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 00:02:14 on December11,2007

    Entry number 219784

    keyword=Swing shift summary

    16:00 Start shift with no beam. Tuning for Hall C in progress.
    20:30 Beam returns. Doing OTR check, Compton rate check (requires Compton DAQ restart), harp scans. All ok,
    21:05 Start production data taking with E_beam = 739 MeV, theta = 120 degrees.
    21:05--24:00 Smooth data taking.

    Run summary
    Left arm:


    Run# p (MeV) Target
    3681 711.973 Fe (right)
    3682 C (right)
    3683 Pb
    3684 663.820 Pb
    3685 C (right)
    3686 Fe (right)
    3687 618.924 Fe (right)
    3688 Fe (left)
    3689 C (left)
    3690 dummy
    3692 Pb
    3693 4He
    3694 577.064 4He
    3695 Pb
    3696 dummy

    Right arm:
    Run# p (MeV) Target
    21726 711.973 Fe (right)
    21727 C (right)
    21728 Pb
    21729 618.924 Fe (right)
    21730 Fe (left)
    21731 C (left)
    21732 dummy
    21734 Pb
    21735 4He